# Maximal Extractable Value in Decentralized Finance: Taxonomy, Detection, and Mitigation

### v1

@@ 9 Challenges and Possible Solutions
Layer-2 MEV: Current detection approaches and mitigation strategies primarily focus on MEV extracted from transactions on the Ethereum Mainnet, a Layer 1 solution. Layer 2 solutions, such as Arbitrum and Optimism, are built on top of Layer 1 to enhance scalability, increase transaction throughput, and reduce costs. Searchers may exploit transactions on Layer 2 to extract MEV, either by performing cross-layer transactions (involving both Layer 1 and Layer 2) or by targeting only Layer 2 transactions. Sjursen et al. [56] conducted a preliminary identification of arbitrage transactions across four networks: Ethereum Mainnet, Arbitrum, Optimism, and Polygon PoS. However, their study was based on manual transaction analysis. There is a growing need for intelligent, cross-layer detection and mitigation to address MEV exploitation across both Layer 1 and Layer 2 networks.
@@
Comprehensive MEV detection: Current MEV detection approaches fail to capture all transaction categories within each MEV type, as shown in Table III. However, developing a comprehensive detection method is challenging, especially as searchers continuously innovate in the the evolving DeFi landscape. Consequently, there is a pressing need to develop a more generalized detection approach that minimizes false negatives. This can be achieved by critically analyzing MEV transaction patterns and detecting MEV based on profits rather than solely on swap patterns.
@@
10 Conclusion
@@
Our findings revealed that challenges such as centralization risks, multi-address MEV, accurate detection, and vulnerabilities in layer-2 solutions remain unresolved despite significant advancements. Addressing these challenges is crucial to fostering a more secure, decentralized, and efficient DeFi environment. Further research is still needed to refine detection approaches, enhance mitigation strategies, and explore new solutions to democratize MEV. Additionally, proactive detection approaches are required to identify MEV transactions in real time and prevent their execution. This comprehensive survey will guide researchers, developers, policymakers, and stakeholders in analyzing and understanding the MEV landscape, offering insights to curb the negative impacts of MEV while unlocking its potential benefits for the DeFi ecosystem.

### v2

@@ 9 Challenges and Possible Solutions
Cross-domain MEV: Current detection approaches and mitigation strategies primarily focus on MEV extracted from transactions on the Ethereum Mainnet, a Layer 1 solution. However, the detection of cross-domain MEV, which consists of DEX-CEX, cross-chain, cross-layer, and cross-rollup, remains largely unaddressed. [32] has explored DEX-CEX arbitrage detection, while [33] focuses on cross-chain arbitrage. Moreover, [70, 34] has provided insights into the manual analysis of cross-layer and cross-rollup arbitrage transactions. There is a pressing need for solutions that can proactively detect and mitigate MEV exploitation across multiple domains. One approach for mitigating cross-rollup MEV is the use of sequencers, which control the ordering of transactions within rollups. Mechanisms like transaction ordering and batch auctions (Section 7.1) can prevent sequencers from manipulating transaction ordering for MEV exploitation. However, mitigating cross-chain MEV is more complex. In particular, bridging protocols, which are used to transfer tokens and assets between different blockchains, introduce delays, creating new opportunities for cross-chain MEV exploitation. Consequently, it becomes crucial to develop secure and efficient cross-domain communication mechanisms and bridging protocols for mitigating MEV associated with cross-chain, cross-layer, and cross-rollup.
@@
Comprehensive MEV detection: Current MEV detection approaches fail to capture all transaction categories within each MEV type, as shown in Table II. Even when these detection approaches are integrated, some MEV transactions remain undetected due to transaction structural complexity and the rigid heuristic rules of these approaches. Developing a comprehensive detection method is challenging, especially as searchers continuously innovate in the evolving DeFi landscape. Consequently, there is a pressing need to develop a more generalized detection approach that minimizes false negatives. This can be achieved by critically analyzing MEV transaction patterns and detecting MEV based on profits rather than solely on swap patterns.
@@
Proactive MEV detection: Most existing MEV detection approaches identify MEV transactions only after they are included in the blockchain (i.e., reactive). Proactive detection, such as pre-chain transaction screening approaches [37, 24], aim to enhance the security and fairness of the DeFi ecosystem by identifying MEV transactions before they impact the network. These approaches function as early-warning systems that flag MEV transactions, enabling proactive countermeasures. However, several challenges remain in proactive detection. One major issue is feature extraction complexity, as distinguishing MEV transactions from non-MEV requires identifying dynamic and evolving MEV patterns, particularly in differentiating value-creating from value-diverting MEV transactions. Furthermore, models trained on historical data may quickly become obsolete due to the rapid evolution of MEV strategies, necessitating adaptive learning approaches. Proactive detection may also face scalability challenges, as real-time analysis requires significant computation resources, especially when integrating learning approaches. Lastly, submission of transactions via privacy-preserving public pools or private pools enables MEV transactions to bypass detection approaches, reducing the effectiveness of pre-chain screening.
Simulation environment: MEV extraction and optimization methods are often evaluated in simulated environments or local forked chains, such as Ganache. However, ensuring that these environments accurately represent real-world conditions, such as network latency, price slippage, and mempool behavior, remains a challenge. Financial gains observed in simulations may not align with those in actual networks, and, in the worst case, a searcher could incur losses in real-world scenarios. For instance, a victim transaction that has already been included in a block might still appear in the mempool to a searcher due to network latency. The searcher’s MEV transactions based on this victim might not yield the desired financial gains. Consequently, it is critical to model network latency and slippage in MEV simulations. Furthermore, simulating adversarial conditions, where multiple searchers compete to extract and optimize the same MEV opportunities, is essential for more accurate quantification of real-world gains.
10 Conclusion and Future Work
@@
Our findings revealed that challenges such as centralization risks, multi-address MEV, accurate detection, and vulnerabilities in layer-2 solutions remain unresolved despite significant advancements. Addressing these challenges is crucial to fostering a more secure, decentralized, and efficient DeFi environment. Further research is still needed to refine detection approaches, enhance mitigation strategies, and explore new solutions to democratize MEV. Additionally, proactive detection approaches are required to identify MEV transactions in real-time and prevent their execution. This comprehensive survey will guide researchers, developers, policymakers, and stakeholders in analyzing and understanding the MEV landscape, offering insights to curb the negative impacts of MEV while unlocking its potential benefits for the DeFi ecosystem.
While our survey provides a qualitative analysis of various MEV detection approaches and mitigation strategies, future work will focus on incorporating quantitative evaluations for a more detailed and objective assessment. This will help guide practitioners and policymakers in selecting the most effective strategies based on specific use cases.